# Differential graded division algebras, their modules, and dg-simple algebras

### v1

@@ Abstract.
We shall dive the definition of a dg-division algebra, that is a concept of a differential graded algebra which may serve as an analogue of a division algebra. Using this concept, we study their dg-modules, give examples when they occur and study their relation to dg-simple algebras.
@@ Introduction
In the present paper we first define a dg-division algebra as a dg-ring without non trivial left or right dg-ideals. The question of an appropriate concept for a dg-division algebra was posed recently by Violeta Borges Marques and Julie Symons. We show that left dg-Noetherian and dg-artinian dg-division algebras are precisely those for which the ring of cycles is a ${\mathbb{Z}}$-graded-division algebras (cf [12, page 38] for the graded concept). For this result we need a technical assumption, namely that the set of left regular elements of the ring of cycles coincides with the set of right regular elements. This holds true, by a result of Goodearl and Stafford [9] for graded-Noetherian and graded-prime rings of cycles, using a graded version of Goldie’s result. We further show that dg-modules over dg-division algebras are free. Not surprisingly, we prove that endomorphism complexes of dg-simple dg-modules over dg-algebras are dg-division algebras. We provide further examples.
@@ 2. Relating dg-division algebras with gr-division algebra
Let $(A,d)$ be a dg-division algebra, and suppose that the set of homogeneous left regular elements of $\ker(d)$ coincides with the set of homogeneous right regular elements of $\ker(d)$. Then $H(A,d)$ is a gr-division algebra.
@@
Let $(A,d)$ be a dg-simple dg-algebra. If $(A,d)$ is left dg-artinian. Then, up to isomorphism of dg-modules, there is a unique simple dg-module $(S,\delta)$ over $(A,d)$.
@@ 3. Dg-modules over Dg-division algebras
and $A\cdot m\simeq\ _{A}A$, i.e. each direct factor is isomorphic to the regular module.
@@
$A$ $\stackrel{{\scriptstyle\rho_{m}}}{{\longrightarrow}}$ $A\cdot m$
@@ 4. The graded center
Remark 4.6.
Recall from [16, Lemma 4.22] that under the hypotheses and in the notations of Theorem 4.5 and up to isomorphism there is a unique dg-simple dg-module $L_{D_{A}D_{B}}$ over $D_{A}\otimes_{Z}D_{B}$.
@@ 5. Dg-Primitive dg-algebras
Corollary 5.7.
Let $(A,d)$ be a dg-algebra and suppose that $(A,d)$ is left dg-artinian and left dg-Noetherian. Then
[…]
Proof. Since $(A,d)$ is dg-artinian and dg-Noetherian it allows a dg-composition series, and hence admits a dg-simple module $(S,\delta)$. By Lemma 5.6 if $(A,d)$ is dg-simple. we get that $(A,d)$ is dg-simple, and if $(A,d)$ is assumed to be left dg-primitive, then $(A,d)$ is dg-prime. This shows the two implications $\Leftarrow$’.
Suppose now that $(A,d)$ is dg-prime. Since $(A,d)$ is supposed to be left dg-artinian and left dg-Noetherian, we get that $dgrad_{2}(A,d)$ is nilpotent by the dg-Nakayama Lemma (cf [16]). Since $(A,d)$ is dg-prime, we get that $dgrad_{2}(A,d)=0$. By [16, Proposition 4.28] we get that $(A,d)$ is dg-simple. This prove that the left most statement implies the right most statement, and with the first step the equivalence of all three.
@@ 6. A differential graded Jacobson-Chevalley density theorem
Let $(A,d)$ be a differential graded algebra and suppose that $(A,d)$ is left dg-Noetherian and left dg-artinian. By Corollary 5.7 we get that $(A,d)$ is dg-simple if and only if $(A,d)$ is dg-primitive. Hence there is a faithful dg-simple dg-module $(S_{(A,d)},\delta)$ over $(A,d)$. Since $(A,d)$ is dg-simple, [16, Proposition 4.28] shows that any other dg-simple dg-module over $(A,d)$ is isomorphic to $(S_{(A,d)},\delta)$. By Proposition 2.9 we get that $End^{\bullet}_{A}(S_{(A,d)},\delta)$ is a dg-division algebra.

### v2

@@ Abstract.
We give the definition of a dg-division algebra, that is a concept of a differential graded algebra which may serve as an analogue of a division algebra. Using this concept, we study their dg-modules, give examples when they occur and study their relation to dg-simple algebras.
@@ Introduction
In the present paper we first define a dg-division algebra as a dg-ring without non trivial left or right dg-ideals. The question of an appropriate concept for a dg-division algebra was posed recently by Violeta Borges Marques and Julie Symons. We show that under a certain mild technical condition dg-division algebras are precisely those for which the ring of cycles is a ${\mathbb{Z}}$-graded-division algebras (cf [12, page 38] for the graded concept). For this result we need a technical assumption, namely that the set of left regular elements of the ring of cycles coincides with the set of right regular elements. This holds true, by a result of Goodearl and Stafford [9] for graded-Noetherian and graded-prime rings of cycles, using a graded version of Goldie’s result. We further show that dg-modules over dg-division algebras are free. We then classify completely dg-division rings under these conditions. Namely, they are either acyclic, and hence the structure is completely given by a result from Aldrich and Garcia Rozas [1] and Nastacescu and van Oystaen [12, A.1.4.3 Corollary] or else they have differential 00. Not surprisingly, we prove that endomorphism complexes of dg-simple dg-modules over dg-algebras are dg-division algebras. We provide further examples.
@@ 2. Relating dg-division algebras with gr-division algebra
Theorem 2.8.
Let $(A,d)$ be a dg-division algebra and suppose that the set of homogeneous left regular elements of $\ker(d)$ coincides with the set of homogeneous right regular elements of $\ker(d)$. Then either $(A,d)$ is acyclic or $d=0$. Further, $\ker(d)=:R$ is either concentrated in degree 00, or else $\ker(d)$ is a twisted Laurent polynomial ring $\ker(d)=R_{0}[T,T^{-1};\sigma]$ for some automorphism $\sigma$ of $R_{0}$ with $Xr=\sigma(r)X$ for all $r\in R_{0}$.
Proof. Indeed, since $H(A,d)$ is a dg-algebra and since $\ker(d)/\textup{im}(d)=H(A,d)$ as anaalgebra, $\textup{im}(d)$ is a twosided ideal of $\ker(d)$. However, as $\ker(d)$ is a gr-division algebra, every homogeneous non zero element of $\ker(d)$ is invertible in $\ker(d)$. Therefore, either $\ker(d)=\textup{im}(d)$, or $\textup{im}(d)$ does not contain any non zero homogeneous element of $\ker(d)$. Hence, either $H(A,d)=0$ in the first case, or in the second case $d=0$. The structure of $\ker(d)$ follows from Nastacescu-van Oystaen [12, A.1.4.3 Corollary].
Example 2.9.
Both cases mentioned in Theorem 2.8 do occur. Let $A=K[X,X^{-1}]$ the Laurent polynomial ring over a field $K$. Then the differential 00 yields a gr-field, whence also a dg-field. However, also $d(X^{2n+1}):=X^{2n}$ and $d(X^{2n})=0$ for all $n\in{\mathbb{N}}$ yields a dg-algebra, which is acyclic, and a dg-field, since it is already a gr-field. But also the cycles is a Laurent polynomial algebra over $K$ in all even degrees, which is again a gr-field.
@@
Let $(A,d)$ be a dg-division algebra, and suppose that the set of homogeneous left regular elements of $\ker(d)$ coincides with the set of homogeneous right regular elements of $\ker(d)$. Then, either $(A,d)$ is acyclic, or $\ker(d)=H(A,d)$ is a gr-division algebra.
@@
Further, we recall
Lemma 2.11.
Remark 2.12.
Note that using Lemma 2.11, if $(A,d)$ is acyclic, then [1, Proposition 4.3] gives more explicitly $A\simeq\ker(d)[X;D]/(X^{2}-y^{2})$, with $d(y)=1$ and $D(a)=-(-1)^{|a|}d(yay)$ for any homogeneous $a\in\ker(d)$. Furthermore, $A=\ker(d)\oplus\ker(d)y$, and the isomorphism is $\Phi:\ker(d)[X;D]\longrightarrow A$ by $\Phi(X)=y$ and $\Phi(a)=a$ for any $a\in\ker(d)$. For any homogeneous $a,b\in\ker(d)$ we get $d(b+ay)=a$. Moreover, $D(a)=Xa-(-1)^{|a|}aX$ for any homogeneous $a\in\ker(d)$. Recall that a dg-division algebra has either differential 00 or $(A,d)$ is acyclic. The cycles are still given by $\ker(d)$, which need to be a gr-division ring.
@@
Let $(A,d)$ be a dg-simple dg-algebra. If $(A,d)$ is left dg-artinian. Then, up to isomorphism of dg-modules and shift in degrees, there is a unique simple dg-module $(S,\delta)$ over $(A,d)$.
@@ 3. Dg-modules over Dg-division algebras
and $A\cdot m\simeq\ _{A}A[-|m|]$, i.e. each direct factor is isomorphic to the regular module.
@@
$A[-|m|]$ $\stackrel{{\scriptstyle\rho_{m}}}{{\longrightarrow}}$ $A\cdot m$
@@ 6. A differential graded Jacobson-Chevalley density theorem
Let $(A,d)$ be a differential graded dg-simple algebra and suppose that $(A,d)$ is left dg-Noetherian and left dg-artinian. By Lemma 5.6 $(A,d)$ is dg-primitive. Hence there is a faithful dg-simple dg-module $(S_{(A,d)},\delta)$ over $(A,d)$. Since $(A,d)$ is dg-simple, [16, Proposition 4.28] shows that any other dg-simple dg-module over $(A,d)$ is isomorphic to $(S_{(A,d)},\delta)$. By Proposition 2.13 we get that $End^{\bullet}_{A}(S_{(A,d)},\delta)$ is a dg-division algebra.